A package of seeds states that the probability of germination is 0.92. For a package containing 75 seeds, the mean and variance of the number of seeds germinating are?
Answer by jim_thompson5910(28717)

Mean = np = 75*0.92 = 69

Variance = np(1-p) = 75*(0.92)*(1-0.92) = 5.52